# Paint Outside the Box: Synthesizing and Selecting Training Data for Visual Grounding

### v1

@@ 4.6 Model Architectures
Ours 31.37 21.25 29.98 27.53+5.79
Dynamic-MDTRE [60] (#Paras = 160M)
@@
Ours 33.55 23.93 35.03 30.84+7.03
QRNet [90] (#Paras = 273M)
@@
Ours 37.29 26.65 35.73 33.22+4.41
5 Related Work
Visual Grounding. Early approaches addressed visual grounding in a two-stage framework [39, 93, 41], which heavily depended on the quality of proposals and large computational costs due to the large number of proposals. To overcome the limitations, more efficient one-stage methods have been developed [11, 29, 90, 63, 12, 76, 60, 77, 82]. TransVG [11] incorporates a DETR encoder for visual feature extraction and introduces the first transformer-based framework for visual grounding. QRNet [90], VG-LAW [63], and TransVG++ [12], integrate language-guided knowledge within the intermediate layers of visual backbones to compute query-dependent visual attention dynamically. As the first diffusion-based framework, LG-DVG [9] progressively reasons the queried object box by denoising a set of noisy boxes with the language guide. Recent works [21, 86] employ GradCAM [58] within the ALBEF model [35] to produce gradient-based visual explanations, represented as heatmaps for grounding outputs. Alternatively, other works [5, 6, 97] propose to train grounding multimodal large language models on extensive fine-grained labeled datasets. Previous works typically train the model on large-scale visual grounding datasets. In contrast, we focus on investigating the feasibility of effectively learning visual grounding under data-scarce settings.
Training with Synthetic Data. Building off image generation and image captioning, training on synthetic data has emerged as a promising research direction [51, 32, 87, 57, 80, 67, 28, 23, 70, 13]. In language tasks, research examines the generation of additional textual training data [17, 25, 48, 49, 72, 74, 89, 75, 56] or visual data conditioned on the text [65, 42, 19, 98, 44, 66, 85, 43] for data augmentation. In computer vision, [3, 91, 69] have demonstrated the benefits of augmenting real images with synthetic images as training data for image classification. In fine-grained understanding tasks, researchers also explore the use of synthetic data in object detection [15, 73, 7, 95, 16, 64], image segmentation [83, 38, 78, 14, 96], and visual grounding [22, 71]. [38, 95, 16] introduce a task-specific head into the image generator, leveraging intermediate-layer features of the generator. Once trained with the under the supervision of an off-the-shelf teacher model, the generator can synthesize images with annotation. In addition, some works [96, 14, 22] first generate images and then employ an auxiliary model to annotate the synthetic data. For instance, DiverGen [14] generate images by a vanilla diffusion model and subsequently depends on SAM [31] for dense annotation. Other methods [15, 7, 83, 73] employ visual priors (e.g., segmentation masks or edges) to guide the synthetic image generation process via controllable diffusion models. For example, FreeMask [83] uses FreestyleNet [81] to synthesize training images conditioned on target dataset-provided semantic masks. However, these methods rely on either the data generator, the annotator, or both to be pretrained on large-scale densely annotated datasets. In contrast, we leverages generative models trained solely on image-caption pairs, which make our method more flexible.
Data Selection. Recent advancements in data selection methods have gained significant attention from the research community [79, 27, 45]. These methods typically employ pre-defined criteria to assign a scalar score to each training example. While simple heuristics like repetition [33], string existence [53], or semantic similarity [1] are commonly used, more sophisticated criteria have also been explored, including gradient norms [52], forgetfulness [68], and influence function scores [84]. A common strategy is to prioritize data samples with lower training loss for subsequent training [88, 59, 8]. This approach leverages the observation that DNNs initially learn generalizable patterns before gradually overfitting to noisy data [2, 61, 37]. To counter confirmation bias, techniques such as co-training models [46, 20, 94, 34] have been proposed, where two networks are trained concurrently and use their combined predictions for selection. MentorNet [26] further extends this idea by introducing a learned curriculum to guide the selection of training data for a student network. In contrast to prior work, we introduce a novel filtering scheme tailored to our data-scarce setting. It select the most effective synthetic data from the hardness and overfitting perspectives.
6 Conclusions
In this paper, we introduce the POBF framework to effectively train a visual grounding model in challenging data-scarce scenarios. By leveraging advanced generative models pretrained solely on image-caption pairs, we synthesize training data efficiently. Specifically, we address the label misalignment issue by inpainting outside the bounding box. Additionally, to identify the most effective synthetic training samples, we propose a filtering scheme comprising three components: a hardness score, an overfitting score, and a penalty term. Experimental results show that our framework achieves an average improvement of 5.83% across four datasets, outperforming leading baselines with an accuracy increase ranging from 2.74% to 4.35%. Ablation studies highlight that our inpainting strategy and filtering scheme provide substantial accuracy gains of 3.53% and 1.96%, respectively. We further validate the robustness and generalizability of POBF across various generative models, data ratios, and model architectures. Finally, our framework sets a new direction for visual grounding in data-scarce settings, offering a scalable and efficient solution for real-world applications where annotated data is limited.

### v2

@@ 5.6. Model Architectures
POBF (Ours) 31.37 21.25 29.98 27.53+5.79
Dynamic-MDTRE (Shi et al., 2023) (#Paras = 160M)
@@
POBF (Ours) 33.55 23.93 35.03 30.84+7.03
QRNet (Ye et al., 2022) (#Paras = 273M)
@@
POBF (Ours) 37.29 26.65 35.73 33.22+4.41
5.7. Case Study
Figure 5. Qualitative examples illustrating the effectiveness of the two proposed scores. The green score indicates the synthetic image with the higher corresponding score.
Generative models may occasionally produce ambiguous examples that degrade model performance. They can also generate synthetic images that closely resemble real ones, increasing the risk of overfitting. In Figure 5, we present two representative examples illustrating how the hardness score ($S_{1}$) and the overfitting score ($S_{2}$) help to mitigate these two issues and select the effective data.
In Figure 5(a), the second synthetic image contains a small cat in the background, which may confuse the model when attempting to localize the target object small cat. Since this example is difficult for the teacher model to localize correctly, it receives a low $S_{1}$ and could thus be filtered out.
In Figure 5(b), the background of the second generated image is visually very similar to the real image — the color, position, and posture of the two buses closely resembles those in the real image. Including such a similar sample in training data may significantly increase the risk of overfitting. In contrast, the first generated image background presents a bus with a noticeably different appearance and placement compared to the real image. As a result, the teacher model assigns a higher $S_{2}$ to the first image, indicating it is less likely to induce overfitting and is more suitable for training.
6. Conclusions
In this paper, we introduced POBF, a novel framework for effectively learning visual grounding in data-scarce scenarios. Leveraging advanced generative models pretrained solely on image–caption pairs, POBF enables high-quality synthetic data generation. By avoiding reliance on specialized networks trained on extensive dense annotations, POBF gains enhanced flexibility and generalizability. Specifically, we proposed a new data generation strategy—paint outside the box—which addresses the label misalignment issues commonly encountered in previous approaches. Additionally, to select the most effective synthetic samples, we introduced a novel filtering scheme tailored for inpainted data. This scheme integrates a hardness score, an overfitting score, and a penalty term. Experimental results demonstrate that POBF achieves an average improvement of 5.83% over the real-data-only baseline across four benchmark datasets. It also surpasses leading baselines, with accuracy gains ranging from 2.74% to 4.35%. Ablation studies confirm that our data generation strategy and filtering scheme contribute significant improvements of 3.53% and 1.96%, respectively. We further validated the robustness and generalizability of POBF across different generative models, training data sizes, and model architectures, and illustrated the effectiveness of our filtering scheme through detailed case analysis. Overall, we believe our approach offers valuable insights and opens new avenues for future research in training data generation and data selection from new perspectives.